# Xu Shen  and  Yuqiang Zheng Morningside Center of Mathematics Academy of Mathematics and Systems Science Chinese Academy of Sciences

### v2

@@ 1. Introduction
This paper deals with the mod $p$ geometry and arithmetic of some PEL-type Shimura varieties for ramified groups. More precisely, we study the reduction modulo $p$ of the splitting models constructed by Pappas-Rapoport in [47]. For smooth splitting models, we explore the mod $p$ Hodge structures on their special fibers by constructing universal $F$-zips with additional structure of fixed type (given by the Hodge cocharacter). Then we deduce some consequences on the geometry and coherent cohomology.
@@
PEL moduli spaces are central objects to study in arithmetic geometry and Langlands program. If the associated reductive group is unramified at $p$, Kottwitz constructed smooth integral PEL moduli spaces in [28]. In the general setting, Rapoport and Zink in [53] introduced similar and generalized integral PEL moduli spaces $\operatorname{\mathscr{A}}^{\operatorname{naive}}$. These are called naive integral models, as Pappas proved in [43] that in some ramified unitary case, the moduli scheme $\operatorname{\mathscr{A}}^{\operatorname{naive}}$ fails to be flat over $\operatorname{\mathcal{O}}_{E}$, the ring of integers of the local reflex field $E$. To study integral models of PEL-type Shimura varieties in ramified case, in [47] Pappas and Rapoport introduced the so-called splitting integral models $\operatorname{\mathscr{A}}^{\mathrm{spl}}$. Roughly speaking, in the setting of [47], the associated $p$-adic reductive group $G$ has the form as a Weil restriction and the ramification mainly comes from restriction of scalars. After fixing the data, the model $\operatorname{\mathscr{A}}^{\mathrm{spl}}$ is proposed as a relative moduli space over the naive integral model $\operatorname{\mathscr{A}}^{\operatorname{naive}}$ constructed in [53]. For any scheme $S$ over $\operatorname{\mathcal{O}}_{E}$, $\operatorname{\mathscr{A}}^{\operatorname{naive}}(S)$ classifies abelian schemes with PEL structure $\underline{A}=(A,\lambda,\iota,\alpha)$ over $S$. Fix a sufficiently large field extension $F|E$. For a scheme $S$ over $\operatorname{\mathcal{O}}_{F}$, $\operatorname{\mathscr{A}}^{\mathrm{spl}}(S)$ classifies $(\underline{A},\underline{\mathscr{F}_{\bullet}})$ where $\underline{A}\in\operatorname{\mathscr{A}}^{\operatorname{naive}}(S)$ and $\underline{\mathscr{F}_{\bullet}}$ is a filtration on the cotangent bundle $\omega_{A/S}$, satisfying certain conditions related to the ramification data. Pappas and Rapoport proved that the model $\operatorname{\mathscr{A}}^{\mathrm{spl}}$ over $\operatorname{\mathcal{O}}_{F}$ admits nice properties. In particular, there is a scheme $\mathbb{M}^{\mathrm{spl}}$ (called the splitting local model) over $\operatorname{\mathcal{O}}_{F}$, so that $\operatorname{\mathscr{A}}^{\mathrm{spl}}$ and $\mathbb{M}^{\mathrm{spl}}$ sit into a local model diagram. Moreover, by construction $\mathbb{M}^{\mathrm{spl}}$ can be realized as a twisted product of unramified local models. There is a natural morphism
$\operatorname{\mathscr{A}}^{\mathrm{spl}}\rightarrow\operatorname{\mathscr{A}}^{\operatorname{naive}}.$
Pappas and Rapoport defined the canonical model $\operatorname{\mathscr{A}}$ as the scheme theoretic image of this morphism. Then $\operatorname{\mathscr{A}}$ is a flat integral model over $\operatorname{\mathcal{O}}_{E}$. In case that the group $G$ is tamely ramified and $p\nmid|\pi_{1}(G_{der})|$, then up to Hasse principle the scheme $\operatorname{\mathscr{A}}$ should coincide with the Kisin-Pappas integral models [26]. We refer to [47] for more information and details on the general theory of splitting models, and to [44, 45, 49] for some recent progress on the canonical models.
@@
The most well known splitting models come from the examples of Hilbert modular varieties. In this case, $\operatorname{\mathscr{A}}=\operatorname{\mathscr{A}}^{\operatorname{naive}}$ and the special fiber of $\operatorname{\mathscr{A}}$ was studied previously by Deligne and Pappas in [11]. Let $L$ be a totally real field of degree $g>1$ and $p$ a prime number. Let $\kappa|\mathbb{F}_{p}$ be a large enough finite field and $\mathcal{M}^{\textrm{DP}}=\operatorname{\mathscr{A}}\otimes\kappa$ the Deligne-Pappas moduli space over $\kappa$, which parametrizes abelian schemes with real multiplication given by $\operatorname{\mathcal{O}}_{L}$ together with polarization and level structure. If $p$ is unramified in $L$, this is a smooth scheme (a special case of the Kottwitz models [28]). Here we are mainly concerned with the case $p$ ramifies in $L$. Then $\mathcal{M}^{\textrm{DP}}$ is only a normal scheme which is singular (cf. [11]). By contrast, the special fiber of $\operatorname{\mathscr{A}}^{\mathrm{spl}}$ over $\kappa$, in this case denoted by $\mathcal{M}^{\textrm{PR}}$, is smooth and the natural morphism
@@
In the more general PEL setting, recently Bijakowski and Hernandez in [4] studied some aspects of the mod $p$ geometry of splitting models. More precisely, they proved that $\operatorname{\mathscr{A}}^{\mathrm{spl}}$ with maximal level at $p$ is smooth under some conditions on the PEL datum. Roughly, these conditions are to ensure that at a $p$-adic place $v_{i}$, the group has the form $\operatorname{Res}_{F_{i}|\operatorname{\mathbb{Q}}_{p}}H_{i}$ where $H_{i}$ is unramified over $F_{i}$, and the level at $p$ is hyperspecial for these $H_{i}$. Bijakowski and Hernandez also proved the $\mu$-ordinary locus (defined as the maximal Newton stratum) is open and dense in the special fiber $\operatorname{\mathscr{A}}^{\mathrm{spl}}_{0}$ of $\operatorname{\mathscr{A}}^{\mathrm{spl}}$. For this, they introduced a so-called Hodge stratification and proved that the maximal Hodge stratum (which contains the $\mu$-ordinary locus) is open and dense.
In this paper, we essentially work in the same setting as that in [4]. Our first observation is that one can modify the local model diagram between $\operatorname{\mathscr{A}}^{\mathrm{spl}}$ and $\mathbb{M}^{\mathrm{spl}}$ constructed in [47] to a local model diagram between $\operatorname{\mathscr{A}}^{\mathrm{spl}}$ and $\prod_{i,j.l}\mathbb{M}^{\mathrm{loc}}(\mathcal{G}_{i,j}^{l},\mu_{i,j}^{l})$, a product of unramified local models related to the PEL datum, cf. Proposition A.8. This is not quite surprising, as we already mentioned above that the splitting local model $\mathbb{M}^{\mathrm{spl}}=\mathbb{M}^{\mathrm{spl}}(\mathcal{G},\mu)$ was constructed in [47] as a twisted product of unramified local models $\mathbb{M}^{\mathrm{loc}}(\mathcal{G}_{i,j}^{l},\mu_{i,j}^{l})$. Let $G$ be the reductive group defined by the PEL datum, which we assume to be connected (thus we exclude the type D case), and $\{\mu\}$ the attached geometric conjugacy class of Hodge cocharacters of $G$. In the following we often fix a suitable choice of $\mu$ in this conjugacy class. Recall that $E|\mathbb{Q}_{p}$ is the field of definition of $\{\mu\}$ and $F|E$ is a large enough extension (so that $G$ splits over $F$). The group scheme related to the modified local model diagram is $\mathcal{G}^{\mathrm{spl}}$, a reductive group over $\operatorname{\mathcal{O}}_{F}$, which is the reductive model of the split group $G_{F}$. On the other hand, let $\mathcal{G}$ be the parahoric model of $G$ over $\mathbb{Z}_{p}$ associated to the integral PEL datum. The local model diagram of [47] corresponds to a morphism of algebraic stacks
@@
which is more suited to study the canonical model $\operatorname{\mathscr{A}}$ and its related canonical local model. While the local model diagram here corresponds to a morphism
@@
In fact, we can go much further. Recall that the theory of $F$-zips was introduced and studied by Moonen-Wedhorn in [41] as a candidate of mod $p$ Hodge structure. This notion has been promoted and enlarged by Fontaine-Jannsen [15], Drinfeld [13], and Bhatt-Lurie [7]. Here we work with the generalization of $F$-zips in another direction: the notion of $G$-zips for a reductive group $G$ over a finite field, cf. [51, 52]. Back to splitting models of PEL-type Shimura varieties, we will work with smooth splitting models $\operatorname{\mathscr{A}}^{\mathrm{spl}}$ from now on. Let $\kappa$ be the residue field of $\operatorname{\mathcal{O}}_{F}$ and $\operatorname{\mathscr{A}}^{\mathrm{spl}}_{0}$ the special fiber of $\operatorname{\mathscr{A}}^{\mathrm{spl}}$ over $\kappa$.
Theorem 1.1.
@@
Let us first comment on the related reductive groups appearing here. For $\operatorname{\mathscr{A}}^{\mathrm{spl}}_{0}$, the group $\mathcal{G}^{\mathrm{spl}}_{0}$ replaces $\mathcal{G}_{0}^{\textrm{rdt}}$, the maximal reductive quotient of the special fiber $\mathcal{G}_{0}$ of $\mathcal{G}$. The group $\mathcal{G}_{0}^{\textrm{rdt}}$ is mainly related to the geometry of $\operatorname{\mathscr{A}}_{0}$, the special fiber of the canonical model $\operatorname{\mathscr{A}}$. The difference between $\mathcal{G}^{\mathrm{spl}}_{0}$ and $\mathcal{G}_{0}^{\textrm{rdt}}$ reflects the ramification data. Indeed, at a $p$-adic place $v_{i}$, the local factors of the two groups admit the following description:
@@
Now we briefly explain the construction of the universal $\mathcal{G}^{\mathrm{spl}}_{0}$-zip of type $\mu$. Since a general PEL datum is involved, we illustrate the ideas by working with the example of Hilbert-Siegel case. For the general case, see sections 2 and 3. Then the PEL datum $(B,\ast,V,\left\langle\cdot,\cdot\right\rangle,\operatorname{\mathcal{O}}_{B},\Lambda)$ is of type C and $B=L$ is a totally real field which we assume $[L:\mathbb{Q}]>1$. We further assume that there is only one finite place $v$ of $L$ over $p$ and the local field $F_{1}:=L_{v}$ is ramified over $\mathbb{Q}_{p}$, i.e. the ramification index $e=[F_{1}:F_{1}^{\textrm{ur}}]>1$, where $F_{1}^{\textrm{ur}}$ is the maximal unramified sub extension of $F_{1}$ over $\mathbb{Q}_{p}$. Let $f=[F_{1}^{\textrm{ur}}:\mathbb{Q}_{p}]=[\kappa_{1}:\mathbb{F}_{p}]$, where $\kappa_{1}$ is the residue field of $F_{1}$, so that $[F_{1}:\mathbb{Q}_{p}]=ef$. Let $K^{p}\subset G(\mathbb{A}_{f}^{p})$ be a fixed sufficiently small open compact subgroup. Recall $F|F_{1}$ is a fixed large enough extension with residue field $\kappa$. For any $\kappa$-scheme $S$, let $\underline{A}=(A,\lambda,\iota,\alpha)$ be an $S$-point of $\mathscr{A}^{\operatorname{naive}}=\mathscr{A}^{\operatorname{naive}}_{K^{p}}$. Then we have an exact sequence of locally free sheaves of $\operatorname{\mathcal{O}}_{S}$-modules
$0\to\omega_{A/S}\to H_{\operatorname{dR}}^{1}(A/S)\to\operatorname{Lie}_{A^{\vee}/S}\to 0$
@@
In the unramified PEL and Hodge type cases, this exact sequence, together with the Frobenius and Verschiebung morphisms, defines the $F$-zip with additional structure of type $\mu$ over $\operatorname{\mathscr{A}}_{0}=\operatorname{\mathscr{A}}^{\operatorname{naive}}_{0}=\operatorname{\mathscr{A}}^{\mathrm{spl}}_{0}$, cf. [41, 62] and [67]. However, in our ramified case here, it turns out a posterior that the type of the Hodge filtrations constructed in this way will vary (according to the Kottwitz-Rapoport types on $\operatorname{\mathscr{A}}_{0}$, see later). Here we have to take into account the splitting structures and the resolution of singularities $\operatorname{\mathscr{A}}_{0}^{\mathrm{spl}}\rightarrow\operatorname{\mathscr{A}}_{0}$. The rough idea is to based on our modified local model diagram above and guided by the constructions in [57] subsections 3.3 and 3.4.
@@
More precisely, let $(\underline{A},\underline{\mathscr{F}_{\bullet}})$ be an $S$-valued point of ${\mathscr{A}_{0}}^{\mathrm{spl}}$. Then by definition $\underline{A}=(A,\lambda,\iota,\alpha)\in\mathscr{A}^{\operatorname{naive}}(S)$, ¯=$\underline{\mathscr{F}_{\bullet}}=\{\mathscr{F}_{j}^{l}\}$ is a Pappas-Rapoport filtration of $\operatorname{\mathcal{O}}_{F_{1}}\otimes_{\mathbb{Z}_{p}}\operatorname{\mathcal{O}}_{S}$-module $\omega_{A/S}$. To explain this term, we write $\mathcal{H}=H_{\operatorname{dR}}^{1}(A/S)$, which is an $\operatorname{\mathcal{O}}_{F_{1}^{\textrm{ur}}}\otimes_{\mathbb{Z}_{p}}\operatorname{\mathcal{O}}_{S}$-module, hence it has a decomposition $\mathcal{H}=\bigoplus_{j}\mathcal{H}_{j}$, where $\mathcal{H}_{j}$ is the locally free sub $\operatorname{\mathcal{O}}_{S}$-module of $\mathcal{H}$ where $\operatorname{\mathcal{O}}_{F_{1}^{\textrm{ur}}}$ acts by the fixed embedding $\sigma_{j}:\operatorname{\mathcal{O}}_{F_{1}^{\textrm{ur}}}\rightarrow\operatorname{\mathcal{O}}_{F}$. For each $j$, there is a pairing on $\mathcal{H}_{j}$ induced from the pairing on $\mathcal{H}$ coming from the polarization $\lambda$. Similarly, we have a decomposition $\omega_{A/S}=\bigoplus_{j}\omega_{j}$ with $\omega_{j}\subset\mathcal{H}_{j}$ for each $1\leq j\leq f$. The Pappas-Rapoport filtration is by definition a filtration of locally direct $\operatorname{\mathcal{O}}_{S}$-factors of each $\omega_{j}$.

### v3

@@ 1. Introduction
This paper deals with the mod $p$ geometry and arithmetic of some Shimura varieties at ramified places. More precisely, we study the reduction modulo $p$ of the splitting models for PEL-type Shimura varieties constructed by Pappas-Rapoport in [49]. For smooth splitting models, we explore the mod $p$ Hodge structures on their special fibers by constructing universal $F$-zips with additional structure of fixed type (determined by the Hodge cocharacter). Then we derive some consequences on the geometry and coherent cohomology. In particular, we reprove the main results of Bijakowski-Hernandez [6] by different methods. We also clarify and generalize the work of Reduzzi-Xiao [56] in the ramified Hilbert case.
@@
PEL moduli spaces are central objects to study in arithmetic geometry and Langlands program. If the associated reductive group is unramified at $p$, Kottwitz constructed smooth integral PEL moduli spaces in [30]. In the general setting, Rapoport and Zink in [55] introduced similar and generalized integral PEL moduli spaces $\operatorname{\mathscr{A}}^{\operatorname{naive}}$. These are called naive integral models, as shown by Pappas in [45] that in some ramified unitary case, the moduli scheme $\operatorname{\mathscr{A}}^{\operatorname{naive}}$ fails to be flat over $\operatorname{\mathcal{O}}_{E}$, the ring of integers of the local reflex field $E$. To study integral models of PEL-type Shimura varieties in ramified case, in [49] Pappas and Rapoport introduced the so-called splitting integral models $\operatorname{\mathscr{A}}^{\mathrm{spl}}$. Roughly speaking, in the setting of [49], the associated $p$-adic reductive group $G$ has the form of a Weil restriction and the ramification mainly comes from restriction of scalars. After fixing the data, the model $\operatorname{\mathscr{A}}^{\mathrm{spl}}$ is proposed as a relative moduli space over the naive integral model $\operatorname{\mathscr{A}}^{\operatorname{naive}}$, the latter constructed in [55]. For any scheme $S$ over $\operatorname{\mathcal{O}}_{E}$, $\operatorname{\mathscr{A}}^{\operatorname{naive}}(S)$ classifies abelian schemes with PEL structure $\underline{A}=(A,\lambda,\iota,\alpha)$ over $S$. Fix a sufficiently large field extension $F|E$. For a scheme $S$ over $\operatorname{\mathcal{O}}_{F}$, $\operatorname{\mathscr{A}}^{\mathrm{spl}}(S)$ classifies $(\underline{A},\underline{\mathscr{F}_{\bullet}})$ where $\underline{A}\in\operatorname{\mathscr{A}}^{\operatorname{naive}}(S)$ and $\underline{\mathscr{F}_{\bullet}}$ is a filtration on the cotangent bundle $\omega_{A/S}$, satisfying certain conditions related to the ramification data. Pappas and Rapoport proved that the model $\operatorname{\mathscr{A}}^{\mathrm{spl}}$ over $\operatorname{\mathcal{O}}_{F}$ admits nice properties. In particular, there is a scheme $\mathbb{M}^{\mathrm{spl}}$ (called the splitting local model) over $\operatorname{\mathcal{O}}_{F}$, so that $\operatorname{\mathscr{A}}^{\mathrm{spl}}$ and $\mathbb{M}^{\mathrm{spl}}$ sit in a local model diagram. Moreover, by construction $\mathbb{M}^{\mathrm{spl}}$ can be realized as a twisted product of unramified local models. There is a natural morphism $\operatorname{\mathscr{A}}^{\mathrm{spl}}\rightarrow\operatorname{\mathscr{A}}^{\operatorname{naive}}$, which is the composition of the forgetful morphism $\operatorname{\mathscr{A}}^{\mathrm{spl}}\rightarrow\operatorname{\mathscr{A}}^{\operatorname{naive}}_{\operatorname{\mathcal{O}}_{F}}$ over $\operatorname{\mathcal{O}}_{F}$ and the base change projection map $\operatorname{\mathscr{A}}^{\operatorname{naive}}_{\operatorname{\mathcal{O}}_{F}}\rightarrow\operatorname{\mathscr{A}}^{\operatorname{naive}}$. Pappas and Rapoport defined the canonical model $\operatorname{\mathscr{A}}$ as the scheme theoretic image of this morphism. Then $\operatorname{\mathscr{A}}$ is a flat integral model over $\operatorname{\mathcal{O}}_{E}$. In case that the group $G$ is tamely ramified and $p\nmid|\pi_{1}(G_{der})|$, then up to Hasse principle the scheme $\operatorname{\mathscr{A}}$ should coincide with the Kisin-Pappas integral models [28]. We refer to [49] for more information and details on the general theory of splitting models (so far only available in the PEL type case), and to [46, 47, 51] for some recent progress on the canonical models.
@@
The most well known splitting models come from the example of Hilbert modular varieties. In this case, $\operatorname{\mathscr{A}}=\operatorname{\mathscr{A}}^{\operatorname{naive}}$ and the special fiber of $\operatorname{\mathscr{A}}$ was studied previously by Deligne and Pappas in [13]. Let $L$ be a totally real field of degree $g>1$ and $p$ a prime number. Let $\kappa|\mathbb{F}_{p}$ be a large enough finite field and $\mathcal{M}^{\textrm{DP}}=\operatorname{\mathscr{A}}\otimes\kappa$ the Deligne-Pappas moduli space over $\kappa$, which parametrizes abelian schemes with real multiplication given by $\operatorname{\mathcal{O}}_{L}$ together with polarization and level structure. If $p$ is unramified in $L$, this is a smooth scheme (a special case of the Kottwitz models [30]). Here we are mainly concerned with the case $p$ ramifies in $L$. Then $\mathcal{M}^{\textrm{DP}}$ is only a normal scheme which is singular (cf. [13]). By contrast, the special fiber of $\operatorname{\mathscr{A}}^{\mathrm{spl}}$ over $\kappa$, in this case denoted by $\mathcal{M}^{\textrm{PR}}$, is smooth and the natural morphism
@@
In the more general PEL setting, recently Bijakowski and Hernandez in [6] studied some aspects of the mod $p$ geometry of splitting models. More precisely, they proved that $\operatorname{\mathscr{A}}^{\mathrm{spl}}$ with maximal level at $p$ is smooth under some conditions on the PEL datum. Roughly, these conditions are to ensure that at a $p$-adic place $v_{i}$, the group has the form $\operatorname{Res}_{F_{i}|\operatorname{\mathbb{Q}}_{p}}H_{i}$ where $H_{i}$ is unramified over $F_{i}$, and the level at $p$ is hyperspecial for these $H_{i}$. In particular, this excludes the ramified unitary groups (labeled as type (AR) in [6]) as local factors. Bijakowski and Hernandez also proved that the $\mu$-ordinary locus (defined as the maximal Newton stratum) is open and dense in the special fiber $\operatorname{\mathscr{A}}^{\mathrm{spl}}_{0}$ of $\operatorname{\mathscr{A}}^{\mathrm{spl}}$. For this, they introduced a so-called Hodge stratification and proved that the maximal Hodge stratum (which contains the $\mu$-ordinary locus) is open and dense.
In this paper, we essentially work in the same setting as that in [6]. Our first key observation is that one can modify the local model diagram between $\operatorname{\mathscr{A}}^{\mathrm{spl}}$ and $\mathbb{M}^{\mathrm{spl}}$ constructed in [49] to a local model diagram between $\operatorname{\mathscr{A}}^{\mathrm{spl}}$ and $\prod_{i,j.l}\mathbb{M}^{\mathrm{loc}}(\mathcal{G}_{i,j}^{l},\mu_{i,j}^{l})$, a product of unramified local models related to the PEL datum, cf. Proposition A.8. This is not quite surprising, as we already mentioned above that the splitting local model $\mathbb{M}^{\mathrm{spl}}=\mathbb{M}^{\mathrm{spl}}(\mathcal{G},\mu)$ was constructed in [49] as a twisted product of unramified local models $\mathbb{M}^{\mathrm{loc}}(\mathcal{G}_{i,j}^{l},\mu_{i,j}^{l})$. Let $G$ be the reductive group defined by the PEL datum, which we assume to be connected (thus we exclude the type D case), and $\{\mu\}$ the attached geometric conjugacy class of Hodge cocharacters of $G$. In the following we often fix a suitable choice of $\mu$ in this conjugacy class. Recall that $E|\mathbb{Q}_{p}$ is the field of definition of $\{\mu\}$ and $F|E$ is a large enough extension (so that $G$ splits over $F$). The group scheme related to the modified local model diagram is $\mathcal{G}^{\mathrm{spl}}$, a reductive group over $\operatorname{\mathcal{O}}_{F}$, which is the reductive model of the split group $G_{F}$. On the other hand, let $\mathcal{G}$ be the parahoric model of $G$ over $\mathbb{Z}_{p}$ associated to the integral PEL datum. The local model diagram of [49] corresponds to a morphism of algebraic stacks
@@
which is more suited to study the canonical model $\operatorname{\mathscr{A}}$ and its related canonical local model. On the other hand, the local model diagram here corresponds to a morphism
@@
In fact, the above observation on the modified local model diagram leads us to go much further. Recall that the theory of $F$-zips was introduced and studied by Moonen-Wedhorn in [43] as a candidate of mod $p$ Hodge structure. This notion has been promoted and enlarged by Fontaine-Jannsen [17], Drinfeld [15], and Bhatt-Lurie [9]. Here we work with the generalization of $F$-zips in another direction: the notion of $G$-zips for a reductive group $G$ over a finite field, cf. [53, 54]. Back to splitting models of PEL-type Shimura varieties, we will work with smooth splitting models $\operatorname{\mathscr{A}}^{\mathrm{spl}}$ from now on (in particular, this excludes the ramified unitary group case). Let $\kappa$ be the residue field of $\operatorname{\mathcal{O}}_{F}$ and $\operatorname{\mathscr{A}}^{\mathrm{spl}}_{0}$ the special fiber of $\operatorname{\mathscr{A}}^{\mathrm{spl}}$ over $\kappa$.
Theorem 1.1 (Theorem 3.13).
@@
Let us first comment on the related reductive groups appearing here. For $\operatorname{\mathscr{A}}^{\mathrm{spl}}_{0}$, the group $\mathcal{G}^{\mathrm{spl}}_{0}$ replaces $\mathcal{G}_{0}^{\textrm{rdt}}$, the maximal reductive quotient of the special fiber $\mathcal{G}_{0}$ of $\mathcal{G}$. The group $\mathcal{G}_{0}^{\textrm{rdt}}$ is mainly related to the geometry of $\operatorname{\mathscr{A}}_{0}$, the special fiber of the canonical model $\operatorname{\mathscr{A}}$. Recall that there is a natural morphism $\operatorname{\mathscr{A}}^{\mathrm{spl}}_{0}\rightarrow\operatorname{\mathscr{A}}_{0}$. The difference between $\mathcal{G}^{\mathrm{spl}}_{0}$ and $\mathcal{G}_{0}^{\textrm{rdt}}$ reflects the ramification data. Indeed, at a $p$-adic place $v_{i}$, the local factors of the two groups admit the following description:
@@
Now we briefly explain the construction of the universal $\mathcal{G}^{\mathrm{spl}}_{0}$-zip of type $\mu$. In the unramified PEL case, we have $\operatorname{\mathscr{A}}_{0}=\operatorname{\mathscr{A}}^{\operatorname{naive}}_{0}=\operatorname{\mathscr{A}}^{\mathrm{spl}}_{0}$ and the construction is direct, cf. [43, 65] and [70]. However, in the ramified case, the construction becomes rather indirect and complicated. Since a general PEL datum is involved, we illustrate the ideas by working with the example of Hilbert-Siegel case. For the general case, see sections 2 and 3. Then the PEL datum $(B,\ast,V,\left\langle\cdot,\cdot\right\rangle,\operatorname{\mathcal{O}}_{B},\Lambda)$ is of type C and $B=L$ is a totally real field which we assume $L:\mathbb{Q}]>1$. We further assume that there is only one finite place $v$ of $L$ over $p$ and the local field $F_{1}:=L_{v}$ is ramified over $\mathbb{Q}_{p}$, i.e. the ramification index $e=[F_{1}:F_{1}^{\textrm{ur}}]>1$, where $F_{1}^{\textrm{ur}}$ is the maximal unramified sub-extension of $F_{1}$ over $\mathbb{Q}_{p}$. Let $f=[F_{1}^{\textrm{ur}}:\mathbb{Q}_{p}]=[\kappa_{1}:\mathbb{F}_{p}]$, where $\kappa_{1}$ is the residue field of $F_{1}$, so that $F_{1}:\mathbb{Q}_{p}]=ef$. Let $K^{p}\subset G(\mathbb{A}_{f}^{p})$ be a fixed sufficiently small open compact subgroup. Recall $F|F_{1}$ is a fixed large enough extension with residue field $\kappa$. For any $\kappa$-scheme $S$, let $\underline{A}=(A,\lambda,\iota,\alpha)$ be an $S$-point of $\mathscr{A}^{\operatorname{naive}}=\mathscr{A}^{\operatorname{naive}}_{K^{p}}$. Then we have an exact sequence of locally free sheaves of $\operatorname{\mathcal{O}}_{S}$-modules
$0\to\omega_{A/S}\to H_{\operatorname{dR}}^{1}(A/S)\to\operatorname{Lie}_{A^{\vee}/S}\to 0.$
@@
In the unramified case, this exact sequence, together with the Frobenius and Verschiebung morphisms, defines the $F$-zip with additional structure of type $\mu$ over $\operatorname{\mathscr{A}}^{\mathrm{spl}}_{0}$. However, in our ramified case here, it turns out a posteriori that the type of the Hodge filtrations constructed in this way will vary, which reflects the singularities of $\operatorname{\mathscr{A}}_{0}$ in some sense.
Indeed, we have the Kottwitz-Rapoport stratification on $\operatorname{\mathscr{A}}_{0}$ (see later), and on each stratum by [59] we have a $\mathcal{G}_{0}^{\textrm{rdt}}$-zip of certain type (dependent on this stratum). Recall the natural map $\operatorname{\mathscr{A}}_{0}^{\mathrm{spl}}\rightarrow\operatorname{\mathscr{A}}_{0}$. We can pullback these $\mathcal{G}_{0}^{\textrm{rdt}}$-zips to the corresponding preimages of KR strata in $\operatorname{\mathscr{A}}_{0}^{\mathrm{spl}}$. But these do not give the correct object that we want. Here, to construct the $\mathcal{G}^{\mathrm{spl}}_{0}$-zip of fixed type $\mu$ over $\operatorname{\mathscr{A}}^{\mathrm{spl}}_{0}$, we have to take into account the splitting structures. It is here that our modified local model diagram (over $\kappa$)
$\operatorname{\mathscr{A}}^{\mathrm{spl}}_{0}\rightarrow[\mathbb{M}^{\mathrm{loc}}(\mathcal{G}^{\mathrm{spl}},\mu)_{0}/\mathcal{G}^{\mathrm{spl}}_{\kappa}]$
plays the key role. Under our smoothness assumption, the underlying topological space of $\mathbb{M}^{\mathrm{loc}}(\mathcal{G}^{\mathrm{spl}},\mu)_{0}/\mathcal{G}^{\mathrm{spl}}_{\kappa}]$ has only one point. Then the construction is guided by the constructions in [59] subsections 3.3 and 3.4. On the other hand, we remark that it is very hard to work with the original local model diagram $\operatorname{\mathscr{A}}^{\mathrm{spl}}_{0}\rightarrow[\mathbb{M}^{\mathrm{spl}}(\mathcal{G},\mu)_{0}/\mathcal{G}_{\kappa}]$ of [49]. The quotient stack $\mathbb{M}^{\mathrm{spl}}(\mathcal{G},\mu)_{0}/\mathcal{G}_{\kappa}]$ is quite complicated, and in fact it is not known whether the set of $\operatorname{\mathcal{G}}_{\kappa}$-orbits on $\mathbb{M}^{\mathrm{spl}}(\mathcal{G},\mu)_{0}$ is finite or not in general (cf. [3] section 8 and [4]).
@@
Let us discuss in more detail the construction of the $\mathcal{G}^{\mathrm{spl}}_{0}$-zip of fixed type $\mu$ over $\operatorname{\mathscr{A}}^{\mathrm{spl}}_{0}$. Let $(\underline{A},\underline{\mathscr{F}_{\bullet}})$ be an $S$-valued point of ${\mathscr{A}_{0}}^{\mathrm{spl}}$. Then by definition $\underline{A}=(A,\lambda,\iota,\alpha)\in\mathscr{A}^{\operatorname{naive}}(S)$, $\underline{\mathscr{F}_{\bullet}}=\{\mathscr{F}_{j}^{l}\}$ is a Pappas-Rapoport filtration of $\operatorname{\mathcal{O}}_{F_{1}}\otimes_{\mathbb{Z}_{p}}\operatorname{\mathcal{O}}_{S}$-module $\omega_{A/S}$. To explain this term, we write $\mathcal{H}=H_{\operatorname{dR}}^{1}(A/S)$, which is an $\operatorname{\mathcal{O}}_{F_{1}^{\textrm{ur}}}\otimes_{\mathbb{Z}_{p}}\operatorname{\mathcal{O}}_{S}$-module, hence it has a decomposition
$\mathcal{H}=\bigoplus_{j}\mathcal{H}_{j},$
where $\mathcal{H}_{j}$ is the locally free sub $\operatorname{\mathcal{O}}_{S}$-module of $\mathcal{H}$ where $\operatorname{\mathcal{O}}_{F_{1}^{\textrm{ur}}}$ acts by the fixed embedding $\sigma_{j}:\operatorname{\mathcal{O}}_{F_{1}^{\textrm{ur}}}\rightarrow\operatorname{\mathcal{O}}_{F}$. For each $j$, there is a pairing on $\mathcal{H}_{j}$ induced from the pairing on $\mathcal{H}$ coming from the polarization $\lambda$. Similarly, we have a decomposition $\omega_{A/S}=\bigoplus_{j}\omega_{j}$ with $\omega_{j}\subset\mathcal{H}_{j}$ for each $1\leq j\leq f$. The Pappas-Rapoport filtration is by definition a filtration of locally direct $\operatorname{\mathcal{O}}_{S}$-factors of each $\omega_{j}$.